X 5.3.17 A borrower had a loan of $10,000.00 at 6% compounded annually, with 9 annual payments. Suppose the borrower paid offPlease show the steps to calculate
Answer #1

Loan Amount P = $10000

Interest Rate = r = 6%

Number of payment periods = N = 9 years

Let monthly payments made be X

Hence, the sum of present value of monthly payments must be equal to the value of the loan amount

=> X/(1+r) + X/(1+r)2 +....+ X/(1+r)N = P

=> X[1- (1+r)-N]/r = P

=> X = rP(1+r)N/[(1+r)N-1]

Let the balance principal after p = 5 years be Z

The Present Value of monthly payments and balance principal should be equal to the loan amount

=> X/(1+r) + X/(1+r)2 + ..... X/(1+r)p + Z/(1+r)p = P

=> X[1- (1+r)-p]/r + Z/(1+r)p = P

substituting X = rP(1+r)N/[(1+r)N-1] in the above equation

=> rP(1+r)N/[(1+r)N-1][1- (1+r)-p]/r + Z/(1+r)p = P

=> [(0.06)(10000)(1+0.06)9/[(1+0.06)9-1]]*[1- (1+0.06)-5]/(0.06) + Z/(0.06)5 = 10000

=> 6193.11 + Z/(1+0.06)5 = 10000

=> Z = (10000 - 6193.11)(1+0.06)5 = $5094.47